# NeuroClean: A Generalized Machine-Learning Approach to Neural Time-Series Conditioning

### v1

@@ 1 Introduction
Electroencephalography (EEG) and Local Field Potential (LFP) data are two widely used techniques to record electrical activity from the brain. Either recorded by means of scalp electrodes in the case of EEGs or of surgically implanted electrodes in specific brain areas, the outcome are time-series of brain activity, along with several sources of noise and multiple artifacts, typically recorded to investigate either normal brain function or the contribution of specific areas to cognition or movement. EEGs, furthermore LFP data, are highly dimensional data, as their signals gather aggregate contributions from groups of neurons firing simultaneously with a high sampling rate via multiple channels. Consequently, physiological and non-physiological noise and artifacts are common occurrences when working with EEG and LFP data ([12]). Typical noise sources are electrical wiring, eye blinks, muscle spasms, eye movements, which vary greatly in intensity and statistics. Consequently, systems and schemes have been developed to remove and mitigate these noisy components. However, most of them involve long hours of semi-manual labor for a researcher to mitigate these problems. This task requires domain-specific knowledge, with automatized and flexible pipelines that offer consistent, objective, and standardized data cleaning, with an efficient data workflow for large datasets. Nevertheless, automated pre-processing pipelines raise several hard challenges in finding the most reliable method to distinguish good from bad data. In particular, we may inadvertently remove neural data losing some fine details of brain activity, relevant to further analyses. Similarly, failing at removing noise and artifacts may result in noise biases when fitting models, and poor reproducibility of research investigations across experiments in different laboratories.
@@
Previous work has created various automatic pipeline schemes that remove specific artifacts, or sets of these artifacts. Most of them make use of Independent Component Analysis (ICA), or variations thereof (such as wavelet enhanced ICA, wICA; [15], [10], ), to reject artifactual components. Significant effort has been devoted to design pipelines for artifact rejection and noise mitigation ([10]; [1]; [9], [14]). Furthermore, some of the pipelines require pre-trained labeled models or structures that may not be available with the specifics of each recording setup arrangement (spatial data, electrode placement, etc.).
@@
To minimize the amount of neural data removed, quantifiable signal quality assessment metrics are needed. Previous work has shown several metrics and performance values for the different pipelines ([8]; [10]; [1]), some focused on visual differences and final signals, and others based on hyperparameter values, i.e., Signal-to-Noise Ratio (SNR), probability of noisy components, number of channels rejected, etc. In addition to these, here we introduce the use of a practical machine learning set-up to assess the utility of the pipeline in live analysis scenarios by tracking the performance of different classification models. Furthermore, this kind of ML analysis provides the flexibility of being capable of coping with different datasets recorded from various parts of the brain and targets of interest (language, motor, attention, etc.).
@@ 2.1 The NeuroClean Pipeline
Bandpass filtering is applied in this pipeline to reduce noise around the signal and focus on brain wave data. That occurs mainly in relevant frequency bands around the alpha, beta, and low-gamma bands. Thus, a Butterworth filter is applied from 1Hz to 500Hz, following steps similar to [2]; where it was found that the application of this filter reduces some of the noise produced by different sources. For any signal that has a lower than 500 Hz sampling rate, the upper band of the band-pass filter is ignored.
@@
Then the ZapLine filter is used to remove the power supply noise from the EEG, as well as line noise that is found within the signal. This method was developed in [3] as a simple yet effective way to remove power line noise and its harmonics from the EEG signal. However, the ZapLine filter requires the frequency of the power supply (60Hz or 50Hz) to find the main frequency and its harmonics. A mix of spectral and spatial filtering is used when the filter is applied by means of two matrix decomposition branches where the target frequencies are removed. To be concise, it is assumed that the first branch $\mathbf{X}^{\prime}$ is perfectly devoid of line artifacts, as the frequencies and harmonics of the line frequency are set to zero. On the other hand, the second branch $\mathbf{X}^{\prime\prime}$ is denoised by using a denoising matrix based on the JD/DSS method ([4]). Lastly, the two branches are mixed back together by adding them. As shown in [3], since the ZapLine method is based on a $1/f_{line}$ squared-shaped filter, the data not on the relevant denoised frequencies is barely affected. The filter was applied with the MEEGkit collection implemented for Python 3.8+.
@@
With the preliminary filtering performed, it is necessary to reject broken, artifact-ridden, or otherwise unusable channels. We remove them by setting the corresponding channels to zero. Broken channels are found using standard deviation (SD) features and an iterative algorithm first described in [13]. Firstly, the algorithm starts by calculating the standard deviation of the signal for the $j$-th channel over the entire sample size of $N$, equation 1.
@@
This step is divided into two parts: first, an independent component analysis (FastICA in [11]) is performed to extract the relevant sources of the data; then a modification of the Multiple Artifact Rejection Algorithm (MARA) proposed in [16] is applied. To be precise, the modification of the MARA consists of using feature clustering instead of pre-trained models. This strategy combines the advantages of being a powerful supervised feature-based component rejection algorithm and a clustering method that can be fitted to any EEG cleaning case. In addition, previous work on ICA component rejection algorithms is based on supervised learning of manual classification data bases ([10]; [17]; [16]); this dependencies on other similar data may result on lower performance on unseen data, and in some cases ([10]) require additional information on the EEG recordings that may not be available.
@@
Figure 2: A summary schematic of the ClusterMARA algorithm, a modification on the MARA algorithm presented by [16]
After computing the features, we are left with a matrix of $s$ components by five features that is used to find clusters in this five-dimensional space. To achieve this, we apply Density-Based Spatial Clustering of Applications with Noise (DBSCAN, [6]) to the matrix, this finds the core samples of high density and expands clusters from them. In this DBSCAN method, a parameter representing the maximum intersample distance to be considered part of the same neighborhood was adjusted to a value of 2; another parameter related to the minimum number of samples in a neighborhood was also set to 2.
@@ 2.2 EEG Cleanliness Quality Assessment
After executing the pipeline, a quality assessment on the input and output is performed in a per-step basis, as well as the general performance of the pipeline. We separate the quality assessment methods into two distinct sections: a traditional one with metrics such as Signal-to-Noise ratio, probability of artifacts, similarity to expected power curves, etc. ([10]; [8]); on the other hand, a machine learning pipeline was created to evaluate the empirical performance of the pipeline.
@@ 2.2.1 Traditional Approach
First, the Signal-to-Noise Ratio (SNR) was calculated before and after each processing step to quantify the relative improvement in signal quality. The SNR was calculated as the ratio between the mean power of the cleaned signal and the residual noise component, following the method described in [8]. An increase in SNR across consecutive stages of the pipeline indicates effective attenuation of non-neural noise sources while preserving task-relevant neural oscillations.
Second, we evaluated the artifact probability for independent components identified during the ICA step. For each component, features such as local skewness, spectral slope, and deviation from the canonical $1/f$ spectral shape were used to estimate the likelihood of being artifactual. These probabilities were derived using a thresholding approach inspired by [16], which allows for an interpretable estimation of the proportion of components removed versus retained.
Third, we assessed the similarity of the cleaned power spectral density (PSD) to the expected $1/f$ spectral distribution, a key characteristic of physiological neural activity ([8]). This similarity was quantified using the Pearson correlation coefficient between the empirical PSD and the theoretical $1/f$ model, calculated for each channel and averaged across the dataset. A higher correlation value indicates better preservation of the physiological spectral structure after artifact removal.
@@ 3.1 Experiment Setup
The first dataset used was a reach-to-grasp behavioral task first described in [7]. Local field potentials and behavioral data were recorded from two adult male Macaca mulatta monkeys. These macaques were trained on a custom-made reach-to-grasp behavioral task. Given instruction cues, they had to reach with the left or right hand and apply pressure to the force transducers using a vertical precision grasp. The force transducers were placed above the starting position (home plate) in front of the custom-made primate chair. The force transducers had several grips that were changed between trials; the precision grip, and the power grip. Grasping the grips required a specific pronation of the forearm, the angle of this pronation could be manipulated by adjusting the grip orientation (0°, 45°, 90°, 135°).
@@ 3.2 Dataset Analyses
The data was also segmented into eight different frequency bands ($\theta$ at 4Hz-7Hz, $\alpha$ at 8Hz-15Hz, $\beta$ at 15Hz-30Hz, low-$\gamma$ at 30Hz-70Hz, high-$\gamma$ at 70Hz-100Hz, low-ripple at 100Hz-150Hz, high-ripple at 150Hz-200Hz, and multi-unit at 200Hz-500Hz similar to [5]) and trained the MLR and 1NN models with all the bands and the complete frequencies data. We obtained the best accuracies in the complete frequencies with the complete pipeline at $\sim$0.97 for MLR and $\sim$0.86 for 1NN, and the second highest in the multi-unit band with $\sim$0.89 for MLR and $\sim$0.78 for 1NN. However, the multi-unit band rose from $\sim$0.81 for MLR and $\sim$0.55 for 1NN for the raw data, and from $\sim$0.78 for MLR and $\sim$0.56 for 1NN for the full frequencies for the raw data. We found that the lower frequencies than the multi-unit band were not very useful when classifying the phases (Pre-Reach, Reach, Grasp) of this motor task. Furthermore, we computed the microaveraged One-Vs-Rest ROC-AUC curves for each frequency band and data step of the pipeline, and the ROC-AUC curves show an increase through each step of the pipeline. Similarly to the accuracy, the area under the curve was maximum in the full frequencies data with ¿0.99 after the pipeline and 0.91 before the pipeline, see 3 F bottom right ROC-AUC curves. For the multi-unit frequency band, the ROC-AUC started directly at 0.96 and increased up to 0.98 after preprocessing. The highest precision was obtained in the Grasp state for the full frequency data with NeuroClean applied at almost perfect classification (¿0.999), see 3 D bottom right square on the full band.

### v2

@@ 1 Introduction
Electroencephalography (EEG) and Local Field Potential (LFP) data are two widely used techniques to record electrical activity from the brain. Either recorded by means of scalp electrodes in the case of EEGs or of surgically implanted electrodes in specific brain areas, the outcome are time-series of brain activity, along with several sources of noise and multiple artifacts, typically recorded to investigate either normal brain function or the contribution of specific areas to cognition or movement. EEGs, furthermore LFP data, are highly dimensional data, as their signals gather aggregate contributions from groups of neurons firing simultaneously with a high sampling rate via multiple channels. Consequently, physiological and non-physiological noise and artifacts are common occurrences when working with EEG and LFP data ([undefk]). Typical noise sources are electrical wiring, eye blinks, muscle spasms, eye movements, which vary greatly in intensity and statistics. Consequently, systems and schemes have been developed to remove and mitigate these noisy components. However, most of them involve long hours of semi-manual labor for a researcher to mitigate these problems. This task requires domain-specific knowledge, with automatized and flexible pipelines that offer consistent, objective, and standardized data cleaning, with an efficient data workflow for large datasets. Nevertheless, automated pre-processing pipelines raise several hard challenges in finding the most reliable method to distinguish good from bad data. In particular, we may inadvertently remove neural data losing some fine details of brain activity, relevant to further analyses. Similarly, failing at removing noise and artifacts may result in noise biases when fitting models, and poor reproducibility of research investigations across experiments in different laboratories.
@@
Previous work has created various automatic pipeline schemes that remove specific artifacts, or sets of these artifacts. Most of them make use of Independent Component Analysis (ICA), or variations thereof (such as wavelet enhanced ICA, wICA; [undefn], [undefi], ), to reject artifactual components. Significant effort has been devoted to design pipelines for artifact rejection and noise mitigation ([undefi]; [undef]; [undefh], [undefm]). Furthermore, some of the pipelines require pre-trained labeled models or structures that may not be available with the specifics of each recording setup arrangement (spatial data, electrode placement, etc.).
@@
To minimize the amount of neural data removed, quantifiable signal quality assessment metrics are needed. Previous work has shown several metrics and performance values for the different pipelines ([undefg]; [undefi]; [undef]), some focused on visual differences and final signals, and others based on hyperparameter values, i.e., Signal-to-Noise Ratio (SNR), probability of noisy components, number of channels rejected, etc. In addition to these, here we introduce the use of a practical machine learning set-up to assess the utility of the pipeline in live analysis scenarios by tracking the performance of different classification models. Furthermore, this kind of ML analysis provides the flexibility of being capable of coping with different datasets recorded from various parts of the brain and targets of interest (language, motor, attention, etc.).
@@ 2.1 The NeuroClean Pipeline
Bandpass filtering is applied in this pipeline to reduce noise around the signal and focus on brain wave data. That occurs mainly in relevant frequency bands around the alpha, beta, and low-gamma bands. Thus, a Butterworth filter is applied from 1Hz to 500Hz, following steps similar to [undefa]; where it was found that the application of this filter reduces some of the noise produced by different sources. For any signal that has a lower than 500 Hz sampling rate, the upper band of the band-pass filter is ignored.
@@
Then the ZapLine filter is used to remove the power supply noise from the EEG, as well as line noise that is found within the signal. This method was developed in [undefb] as a simple yet effective way to remove power line noise and its harmonics from the EEG signal. However, the ZapLine filter requires the frequency of the power supply (60Hz or 50Hz) to find the main frequency and its harmonics. A mix of spectral and spatial filtering is used when the filter is applied by means of two matrix decomposition branches where the target frequencies are removed. To be concise, it is assumed that the first branch $\mathbf{X}^{\prime}$ is perfectly devoid of line artifacts, as the frequencies and harmonics of the line frequency are set to zero. On the other hand, the second branch $\mathbf{X}^{\prime\prime}$ is denoised by using a denoising matrix based on the JD/DSS method ([undefc]). Lastly, the two branches are mixed back together by adding them. As shown in [undefb], since the ZapLine method is based on a $1/f_{line}$ squared-shaped filter, the data not on the relevant denoised frequencies is barely affected. The filter was applied with the MEEGkit collection implemented for Python 3.8+.
@@
With the preliminary filtering performed, it is necessary to reject broken, artifact-ridden, or otherwise unusable channels. We remove them by setting the corresponding channels to zero. Broken channels are found using standard deviation (SD) features and an iterative algorithm first described in [undefl]. Firstly, the algorithm starts by calculating the standard deviation of the signal for the $j$-th channel over the entire sample size of $N$, equation 1.
@@
This step is divided into two parts: first, an independent component analysis (FastICA in [undefj]) is performed to extract the relevant sources of the data; then a modification of the Multiple Artifact Rejection Algorithm (MARA) proposed in [undefo] is applied. To be precise, the modification of the MARA consists of using feature clustering instead of pre-trained models. This strategy combines the advantages of being a powerful supervised feature-based component rejection algorithm and a clustering method that can be fitted to any EEG cleaning case. In addition, previous work on ICA component rejection algorithms is based on supervised learning of manual classification data bases ([undefi]; [undefp]; [undefo]); this dependencies on other similar data may result on lower performance on unseen data, and in some cases ([undefi]) require additional information on the EEG recordings that may not be available.
@@
Figure 2: A summary schematic of the ClusterMARA algorithm, a modification on the MARA algorithm presented by [undefo]
After computing the features, we are left with a matrix of $s$ components by five features that is used to find clusters in this five-dimensional space. To achieve this, we apply Density-Based Spatial Clustering of Applications with Noise (DBSCAN, [undefe]) to the matrix, this finds the core samples of high density and expands clusters from them. In this DBSCAN method, a parameter representing the maximum intersample distance to be considered part of the same neighborhood was adjusted to a value of 2; another parameter related to the minimum number of samples in a neighborhood was also set to 2.
@@ 2.2 EEG Cleanliness Quality Assessment
After executing the pipeline, a quality assessment on the input and output is performed in a per-step basis, as well as the general performance of the pipeline. We separate the quality assessment methods into two distinct sections: a traditional one with metrics such as Signal-to-Noise ratio, probability of artifacts, similarity to expected power curves, etc. ([undefi]; [undefg]); on the other hand, a machine learning pipeline was created to evaluate the empirical performance of the pipeline.
@@ 2.2.1 Traditional Approach
First, the Signal-to-Noise Ratio (SNR) was calculated before and after each processing step to quantify the relative improvement in signal quality. The SNR was calculated as the ratio between the mean power of the cleaned signal and the residual noise component, following the method described in [undefg]. An increase in SNR across consecutive stages of the pipeline indicates effective attenuation of non-neural noise sources while preserving task-relevant neural oscillations.
Second, we evaluated the artifact probability for independent components identified during the ICA step. For each component, features such as local skewness, spectral slope, and deviation from the canonical $1/f$ spectral shape were used to estimate the likelihood of being artifactual. These probabilities were derived using a thresholding approach inspired by [undefo], which allows for an interpretable estimation of the proportion of components removed versus retained.
Third, we assessed the similarity of the cleaned power spectral density (PSD) to the expected $1/f$ spectral distribution, a key characteristic of physiological neural activity ([undefg]). This similarity was quantified using the Pearson correlation coefficient between the empirical PSD and the theoretical $1/f$ model, calculated for each channel and averaged across the dataset. A higher correlation value indicates better preservation of the physiological spectral structure after artifact removal.
@@ 3.1 Experiment Setup
The first dataset used was a reach-to-grasp behavioral task first described in [undeff]. Local field potentials and behavioral data were recorded from two adult male Macaca mulatta monkeys. These macaques were trained on a custom-made reach-to-grasp behavioral task. Given instruction cues, they had to reach with the left or right hand and apply pressure to the force transducers using a vertical precision grasp. The force transducers were placed above the starting position (home plate) in front of the custom-made primate chair. The force transducers had several grips that were changed between trials; the precision grip, and the power grip. Grasping the grips required a specific pronation of the forearm, the angle of this pronation could be manipulated by adjusting the grip orientation (0°, 45°, 90°, 135°).
@@ 3.2 Dataset Analyses
The data was also segmented into eight different frequency bands ($\theta$ at 4Hz-7Hz, $\alpha$ at 8Hz-15Hz, $\beta$ at 15Hz-30Hz, low-$\gamma$ at 30Hz-70Hz, high-$\gamma$ at 70Hz-100Hz, low-ripple at 100Hz-150Hz, high-ripple at 150Hz-200Hz, and multi-unit at 200Hz-500Hz similar to [undefd]) and trained the MLR and 1NN models with all the bands and the complete frequencies data. We obtained the best accuracies in the complete frequencies with the complete pipeline at $\sim$0.97 for MLR and $\sim$0.86 for 1NN, and the second highest in the multi-unit band with $\sim$0.89 for MLR and $\sim$0.78 for 1NN. However, the multi-unit band rose from $\sim$0.81 for MLR and $\sim$0.55 for 1NN for the raw data, and from $\sim$0.78 for MLR and $\sim$0.56 for 1NN for the full frequencies for the raw data. We found that the lower frequencies than the multi-unit band were not very useful when classifying the phases (Pre-Reach, Reach, Grasp) of this motor task. Furthermore, we computed the microaveraged One-Vs-Rest ROC-AUC curves for each frequency band and data step of the pipeline, and the ROC-AUC curves show an increase through each step of the pipeline. Similarly to the accuracy, the area under the curve was maximum in the full frequencies data with ¿0.99 after the pipeline and 0.91 before the pipeline, see 3 F bottom right ROC-AUC curves. For the multi-unit frequency band, the ROC-AUC started directly at 0.96 and increased up to 0.98 after preprocessing. The highest precision was obtained in the Grasp state for the full frequency data with NeuroClean applied at almost perfect classification (¿0.999), see 3 D bottom right square on the full band.